# Polle Norbury Georgescu Cummins Goria LondonUK King’s College LondonUK

### v1

@@ 2.1 Dataset
Our analysis uses data from a large-scale longitudinal study of shift workers [7], comprising 1,185 participants monitored over a two-week period. Study procedures were reviewed by an independent research ethics expert (Dr. David Carpenter) working under the auspices of the Association of Research Managers and Administrators (ARMA, https://arma.ac.uk/, involved in formulating national standards for research ethics support and review) and received a favourable ethics opinion on 22nd November 2023. All participants gave written informed consent. Access to the dataset can be made through reasonable requests to the senior author.
During the study, participants completed speech recordings and self-report assessments twice daily on their personal devices, coinciding with the start and end of their work shifts. From the collected metrics, we focus on time since sleep ($S$) as our target variable, as a function of fatigue [18]. Among the various speech tasks completed by participants, we analysed the reading of fourteen different texts from the speech literature, including the Aesop fables, The North Wind and the Sun [19], The Boy who Cried Wolf [20], and the Rainbow Passage [21]. On each day, different texts (but the same for all participants, although translated into the relevant language, i.e. English or Spanish), were assigned before and after each work shift. These standardized reading tasks provide a controlled context where variations in speech primarily reflect paralinguistic features rather than content or speaking style [22].
@@ 3.2 Benchmark Models
Mixed-effects Model: We implement a mixed-effects approach [17], following [7] to incorporate speaker-specific patterns. For each speaker $i$ and time point $t$, we train a separate ridge regression model using all previous observations ${(x^{i}_{j},y^{i}_{j})}_{j<t}$ from that speaker. Unlike our meta-learning models, which can be trained once and then adapted to new speakers, this approach requires full model retraining with each new observation, making it considerably less practical in deployment. The evaluation methodology for this baseline differs from our other models: following [7]’s approach on the same dataset, we use speakers’ past observations as training data and their current observations as test data. As this model serves as a benchmark for comparison with our proposed methods, we deem this methodological difference acceptable.
@@ 3.3 Meta-learning Approaches
Using a GPT2 architecture (15 layers, 8 heads), we frame fatigue prediction as next-token prediction, computing loss only at target positions (the $y$s). To form the sequences and since transformers require uniform input dimensionality, targets are augmented to match feature dimensionality by sampling each dimension from $\mathcal{N}(y^{i}_{j},0.1^{2})$. We use a mean squared error loss for each dimension of the augmented target, which is then averaged to produce the final loss.
@@ 5 Conclusions and Future Work
Some limitations of our approach should be noted. While our models show promising overall performance – particularly given the variability in devices and recording settings – we observe performance variations across demographic groups, with differences in accuracy for female speakers and those under 40. Additionally, our current analysis uses fixed-length speech samples from standardised reading tasks. The same dataset includes other speech activities such as image descriptions and question-answering tasks, offering opportunities to explore more naturalistic speech patterns. Future work could investigate how our approach generalises across these different speech tasks, potentially enabling more flexible real-world applications that don’t rely on specific text reading.

### v2

@@ 2.1 Dataset
Our analysis uses data from a large-scale longitudinal study of shift workers [7], comprising 1,185 participants monitored over a two-week period. Study procedures were reviewed by an independent research ethics expert (Dr. David Carpenter) working under the auspices of the Association of Research Managers and Administrators (ARMA, https://arma.ac.uk/, involved in formulating national standards for research ethics support and review) and received a favourable ethics opinion on 22nd November 2023. All participants gave written informed consent. Access to the dataset can be obtained through reasonable requests to the senior author.
During the study, participants completed speech recordings and self-report assessments twice daily on their personal devices, coinciding with the start and end of their work shifts. From the collected metrics, we focus on time since sleep ($S$) as our target variable, representing a key component of fatigue [18]. Among the various speech tasks completed by participants, we analysed the reading of fourteen different texts from the speech literature, including the Aesop fables, The North Wind and the Sun [19], The Boy who Cried Wolf [20], and the Rainbow Passage [21]. On each day, different texts (but the same for all participants, although translated into the relevant language, i.e. English or Spanish), were assigned before and after each work shift. These standardized reading tasks provide a controlled context where variations in speech primarily reflect paralinguistic features rather than content or speaking style [22].
@@ 3.2 Benchmark Models
Mixed-effects Model: We implement a mixed-effects approach [17], following [7] to incorporate speaker-specific patterns (with fixed effects for age, sex, and language, and random effects for participants). For each speaker $i$ and time point $t$, we train a separate ridge regression model using all previous observations ${(x^{i}_{j},y^{i}_{j})}_{j<t}$ from that speaker. Unlike our meta-learning models, which can be trained once and then adapted to new speakers, this approach requires full model retraining with each new observation, making it considerably less practical in deployment. The evaluation methodology for this baseline differs from our other models: following [7]’s approach on the same dataset, we use speakers’ past observations as training data and their current observations as test data. As this model serves as a benchmark for comparison with our proposed methods, we deem this methodological difference acceptable.
@@ 3.3 Meta-learning Approaches
Using a GPT2 architecture (15 layers, 8 heads), we frame fatigue prediction as next-token prediction, computing loss only at target positions (the $y$s). To form the sequences, targets are augmented to match feature dimensionality by sampling each dimension from $\mathcal{N}(y^{i}_{j},0.1^{2})$. We use a mean squared error loss for each dimension of the augmented target, which is then averaged to produce the final loss.
@@ 5 Conclusions and Future Work
Some limitations of our approach should be noted. While our models show promising overall performance – particularly given the variability in devices and recording settings – we observe performance variations across demographic groups, with differences in accuracy for female speakers and those under 40. Our investigation is also limited to the time since sleep variable, which is only one component of fatigue. Additionally, our current analysis uses fixed-length speech samples from standardised reading tasks. The same dataset includes other speech activities such as image descriptions and question-answering tasks, offering opportunities to explore more naturalistic speech patterns. Future work could investigate how our approach generalises across these different speech tasks, potentially enabling more flexible real-world applications that don’t rely on specific text reading.